# Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels

### v2

@@ A.2 LLaMa2 and QWEN plots
We show 4 bit to 2 bit variable layer-wise quantization for LLaMa2-7b and QWEN-7B in fig. 8 and fig. 9 respectively. All notations are same as in fig. 2. These curves were also generated on 2K evaluation instances from each of the datasets.
@@ A.3 Commonality between layer importance
A.4 Results of 8bit to 4bit quantization on different datasets
We show results of quantizing models lower than 8 bit. Following our proposed methodology, the same techinque can be applied to have the more important layers in 8 bits and the least important ones in 4 bits. While this does still increase the performance that can be fit within a memory requirement, the results are not as major as the ones for the 4-2 bit range, thus we mainly focus on that range.

### v3

@@ A.2 LLaMa2 and QWEN plots
We show 4-bits to 2-bits variable layer-wise quantization for LLaMa2-7b and QWEN-7B in fig. 8 and fig. 9 respectively. All notations are same as in fig. 2. These curves were also generated on 2K evaluation instances from each of the datasets.
@@ A.3 Commonality between layer importance
A.4 Results of 8-bits to 4-bits quantization on different datasets
We show results of quantizing models lower than 8-bits. Following our proposed methodology, the same techinque can be applied to have the more important layers in 8-bits and the least important ones in 4-bits. While this does still increase the performance that can be fit within a memory requirement, the results are not as major as the ones for the 4-2 bit range, thus we mainly focus on that range.